OpenStudy (some.random.cool.kid):

When solving x2 + 10x + 3 = 0 by completing the square, which of the following is a step in the process? (x + 5)2 = 28 (x + 5)2 = 22 (x − 5)2 = 22 (x − 5)2 = 28

OpenStudy (some.random.cool.kid):

@SamsungFanBoy

OpenStudy (some.random.cool.kid):

@mathstudent55 @Michele_Laino

OpenStudy (anonymous):

Complete the square what do you do first?

OpenStudy (some.random.cool.kid):

thats the issue idk...

OpenStudy (anonymous):

aX^2+bx +c= 0 Step one get rid of the c...So in your equation we need to get +3 to the other side of the equation..how do we do that?

OpenStudy (some.random.cool.kid):

@nelsonjedi

OpenStudy (some.random.cool.kid):

sorry I had to leave for a minute were having guest over and I needed to help clean the house real quick. @nelsonjedi

OpenStudy (some.random.cool.kid):

@Nnesha can you help?

OpenStudy (some.random.cool.kid):

@mathstudent55 @Michele_Laino @SamsungFanBoy @TheSmartOne @triciaal mind helping anyone?

OpenStudy (anonymous):

...the first step is to add subtract 3 from both sides and you get X^2+10x= -3 the set up the equation to complete the square x^2+10x+ c = -3 + c c= (1/2(b))^2 where b=10 So what would c be?

OpenStudy (some.random.cool.kid):

mmm idk?

OpenStudy (anonymous):

Take 1/2 of 10 then square it...that is your c value..that is how you complete the square by taking 1/2 of b and squaring it then adding it to both sides

OpenStudy (some.random.cool.kid):

a hah so 25?

OpenStudy (some.random.cool.kid):

is this going to be a postive ora negative outcome?

OpenStudy (some.random.cool.kid):

nvm I got it ... thanks

OpenStudy (anonymous):

sorry disconnect...so now you have X^2+10X+25 = -3+25

OpenStudy (some.random.cool.kid):

thank you I passed.

OpenStudy (some.random.cool.kid):

and indeed I faced similar troubles with the OS system too @nelsonjedi

OpenStudy (anonymous):

Very good glad I could help